THE ^{1}Π STATES OF NaCs: SPECTROSCOPY, LIFETIMES, PERMANENT AND TRANSITION DIPOLE MOMENTS.

\underline{R. FERBER}, O. DOCENKO, I. KLINCARE, O. NIKOLAYEVA, M. TAMANIS, J. ZAHAROVA, University of Latvia, Department of Physics, LV-1586, Riga, Latvia; A. PASHOV, Sofia University, Department of Physics, 1164 Sofia, Bulgaria; H. KNÖCKEL, E. TIEMANN, Leibniz Universität Hannover, Inst.f. Quantenoptik, 30167 Hannover, Germany; A. V. STOLYAROV, E. A. PAZYUK, A. ZAITSEVSKII, Moscow State University, Department of Chemistry, Moscow, 119899, Russia.

The NaCs molecule is one of the prospective objects for production of ultracold polar molecules. LIF Fourier transform spectroscopy study is presented for the (1,3)^{1}Π states with 0.03 cm^{−1} resolution. Potential energy curves (PECs) are obtained by the Inverted Perturbation Approach reproducing (3)^{1}Π state^{a} energies for R = 3 to 11 Å. In the B(1)^{1}Π state we accounted for numerous B(1)^{1}Π ∼ c^{3}Σ^{+} perturbations by omitting perturbed levels for the fit to construct the PEC for R = 2.6 to 8.4 Å. The permanent electric dipole moments d and the Λ-splitting were measured^{b} by dc Stark mixing and electric RF-optical double resonance methods yielding d within 5 - 8 D for (3)^{1}Π and d ∼ 1 D for the D(2)^{1}Π (v < 3) state. The radiative lifetimes τ were measured from LIF kinetics as τ = 29 to 21 ns for (3)^{1}Π (v = 3 to 25) and τ = 37 ns for D(2)^{1}Π (v = 0). The measured data are supported by electronic structure calculations for the (1-3)^{1}Π states^{b} by many-body multipartitioning perturbation theory of PECs, permanent and transition dipole moments, as well as angular coupling matrix elements for the lowest singlet states. The predicted d values reproduce their experimental counterparts within the measurement errors. Lifetimes for the (1-3)^{1}Π-states have been calculated in Hund’s ”a” coupling case using the approximate sum rule over the lower vibronic states. The spectra and formation rates of ultracold NaCs in the X^{1}Σ^{+} (v = 0, J = 0) state were simulated for the optical cycle a^{3}Σ^{+} → B(1)^{1}Π ∼ c^{3}Σ^{+} ∼ b^{3}Π → X^{1}Σ^{+}. The Riga team and the Moscow team acknowledge support by NATO SfP 978029 Optical Field Mapping grant, the Hannover team support by the DFG through the SFB 407.

aO. Docenko et al J. Chem. Phys., \underline{124}, 174310, 2006.

bJ. Zaharova et al J. Chem. Phys., \underline{124}, 184318, 2006.